OpenStudy (goformit100):

At a dinner party every two guests used a bowl of rice between them, every three guests used a bowl of dal between them and every four used a bowl of meat between them. There were altogether 65 dishes. How many guests were present at the party ? A. 60 B. 65 C. 90 D. None of these

OpenStudy (goformit100):

@Noemi95

OpenStudy (abb0t):

\[\frac{ x }{ 2 } + \frac{ x }{ 3 } + \frac{ x }{ 4 } = 65x\]

OpenStudy (anonymous):

A.60 Let the number of guests be x. Then, number of bowls of rice=x/2; number of bowls of dal=x/3; number of bowls of meat=x/4 So, x/2+x/3+x/4=65 -> 6x+4x+3x/12=65 -> 13x=65*12 x=65*12/13=60